In a survey of 280 adults over 50, 75% said they were taking vitamin supplements. Find the margin of error for this survey if we want a 99% confidence in our estimate of the percent of adults over 50 who take vitamin supplements.

In a survey of 3300 T.V. viewers, 20% said they watch the network news programs. Find the margin of error for this survey if we want a 95% confidence in our estimate of the percent of T.V. viewers who watch network news programs.

1
Expert's answer
2021-05-07T14:16:30-0400

In a survey of 280 adults over 50, 75% said they were taking vitamin supplements. Find the margin of error for this survey if we want a 99% confidence in our estimate of the percent of adults over 50 who take vitamin supplements.

"\hat{p} = 0.75 \\

n = 280 \\

α = 1 - 0.99 = 0.01"

Crotocal z-value

"z_{\u03b1\/2} = z_{0.01\/2} = z_{0.005} = 2.58"

Standard error of "\hat{p}:"

"SE = \sqrt{\frac{\hat{p} \times (1- \hat{p})}{n}} \\

= \sqrt{\frac{0.75 \times 0.25}{280}} \\

= 0.025877 \\

E = z_{α/2} \times SE \\

= 2.58 \times 0.025877 \\

= 0.066764"

99 % Confidence interval:

0.75 ± 0.066764

(0.683, 0.817)


In a survey of 3300 T.V. viewers, 20% said they watch the network news programs. Find the margin of error for this survey if we want a 95% confidence in our estimate of the percent of T.V. viewers who watch network news programs.

The equation for the margin of error is:

"E = Z( \frac{\hat{p}(1- \hat{p})}{n} )^{0.5}"

z = 1.96 for 95 % confidence

"\hat{p} = \frac{X}{n} = \frac{660}{3300} = 0.20 \\

n = 3300 \\

E = 1.96( \frac{0.2 \times 0.8}{3300} )^{0.5} = 0.0136"